OpenStudy (anonymous):

Choose the slope-intercept equation of the line that passes through the point shown and is perpendicular to the line shown. line through the points 0 comma 6 and negative 2 comma 0 with a point 6 comma 2 not on the line y = one thirdx y = -one thirdx + 4 y = 3x + 20 y = -3x - 16

OpenStudy (anonymous):

@jhonyy9

OpenStudy (anonymous):

I will medal and fan

OpenStudy (texaschic101):

slope (m) = (y2 - y1) / (x2 - x1) (0,6)...x1 = 0 and y1 = 6 (-2,0)...x2 = -2 and y2 = 0 now we sub slope(m) = (0 - 6) / (-2 - 0) slope = -6/-2 slope = 3 perpendicular slope needs the negative reciprocal of that slope. All that means is " flip " the slope and change the sign. So we need the negative reciprocal of 3, which is -1/3...see how I flipped the slope and changed the sign. Now we use y = mx + b slope(m) = -1/3 use either of your points...(0,6)...x = 0 and y = 6 y = mx + b 6 = -1/3(0) + b 6 = b so your perpendicular equation is : y = -1/3x + 6 ...but it does not contain (6,2)...hmmm ....there is only one answer it can be since there is a -1/3 slope and that is : y = -1/3x + 4

OpenStudy (anonymous):

OMG!!! Thank you so much @texaschic101